Explain step by step

Golden Oldies is a small company that burns music CDs for non-technical senior citizens. The table below contains information on the steps required to burn and package a CD. (Their equipment is quite old, hence, the long burn time.) No specific demand numbers are available, however, the cycle time should not exceed 30 minutes in order to meet demand requirements.

  Task   Description   Req'd Time (in minutes)   Predecessor
A Set up equipment 10 None
B Test equipment 12 A
C Select songs 26 None
D Burn CD 15 B, C
E Package CD 6 D

(Missing questions from prior question)

Given your task assignments to workstations, what is the actual number of workstations required?


What is the cycle time (in minutes) for this production line?  

What is the combined idle time (in minutes) for this production line?


What is the overall efficiency for this production line? (Write your answer as a percentage, and display your answer to two decimal places.)
Answer #1

Number of workstations will be 3, because the cycle time should not increase 30 sec. and the task time of 69 can be accommodated in minimum 3 workstations.

Cycle time = 26 sec, which is equal to the time taken by longest task.

Idle time for the production line = (26x3) -69 =9 sec. per cycle.

Efficiency of the production line = 69/3x26 = 0.8846